OpenStudy (loser66):

if delta l = alpha * lo * delta T and delta V = beta * Vo * delta T how does beta relate to alpha when l is length and V is volume of a cube Please, help

OpenStudy (loser66):

is beta = alpha ^3? @ganeshie8

ganeshie8 (ganeshie8):

is Vo = lo^3 ?

OpenStudy (loser66):

yes

ganeshie8 (ganeshie8):

if delta l = alpha * lo * delta T and delta V = beta * Vo * delta T Vo = lo^3 so, beta*deltaT = (alpha*delta)^3 beta = alpha^3*deltaT^2

OpenStudy (loser66):

yes, I think so, too.

ganeshie8 (ganeshie8):

change in L must produce change in volume of (deltaL)^3 hmm... i may not be understanding the question correctly though.. whats deltaT ?

OpenStudy (loser66):

It's from Physics, delta T is the change of temperature

OpenStudy (loser66):

The question is about the expansion coefficient of the object linearly depend on length and volume. (extra question)

OpenStudy (loser66):

However, I think by algebraic way, we can find out the link between them like what you did. Thanks a lot. Master

ganeshie8 (ganeshie8):

lol okay hope the cube stays as cube after expansion also :)